Sams Ltd is specialized in producing and selling washing machines. In 2018, the manufacturing cost per unit included:

£ Direct material 125

Direct labor (20 minutes per unit) 45/hour Variable manufacturing overhead 20 Variable selling expenses 15 Variable administrative expenses 10

Fixed costs for the year ended 31 December 2018 were: £000

Fixed manufacturing 1,100 Fixed selling and distribution 1,450 Fixed administrative 675

The company produced and sold 220,000 units at £225 per unit.

In 2019, management has decided to increase the selling price by 25% and to maintain the same contribution margin ratio as last year. This increase in price is to meet an increase of £1,450,000 in fixed costs in 2019. The company has produced and sold the same quantity in 2019 as last year.

Required

(a) Calculate the break-even point and margin of safety in both units and revenue for the two years, 2018 and 2019, and briefly comment upon the results.




1
Expert's answer
2022-02-28T09:51:01-0500

BEP

2018

BEP in units = F.C/contribution unit

Contribution unit= 225- 90 = 135

3225/135 =23.889 units

3225/135*225 = sh 5375

2019

Contribution unit = 281.25-90= 191.25

FC=3225 +1450=4675000

=4675000/199.25 = 24,444units

=4675/191.25*281.25 = sh 6875000

MOS

2019

=actual sales – sales at BEP

(220000*225) - 5375= 49494625

(220000*281.25) – 6875000 =